# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ C.3 Proof for Theorem 3
[…]
@@
$\begin{split}{\mathbb{E}}[R(T)]&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[dD_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16Kd\lambda_{m}}{\Delta}+d\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[dD_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}Kd(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+d\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880dC_{m}+1440d+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}Kd\log(8/\Delta)\log(240K/\Delta)}{\Delta}+d\log(8/\Delta)\\ &=O\left(dC+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{dK\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
Appendix D Proof of Theorem 4
D.1 Notations
D.2 Lemmas for Proving Theorem 4
Lemma 14
For the BB-BARBAT algorithm time horizon $T$, the number of epochs $M$ is at most $\min\{{\mathcal{M}},\log(T)\}$. In the $m$-th epoch, the selected arm $k_{m}$ must satisfy $\Delta_{k_{m}}^{m-1}=a^{-(m-1)}$.
Proof 13
The length of epoch $m$ is given by $N_{m}=K\lambda_{m}a^{2(m-1)}\geq a^{2(m-1)}$. From the lower bound of $N_{m}$ and $a=\max\{T^{\frac{1}{2({\mathcal{M}}+1)}},2\}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{a^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=a^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq a^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold.
Lemma 15
For any epoch $m$, the length $N_{m}$ satisfies $N_{m}\geq\sum_{k\in[K]}n_{k}^{m}$.
Proof 14
Lemma 16
For epoch $s$ and $m$ with $1\leq s\leq m$, the following inequality holds:
$\frac{\lambda_{m}}{\lambda_{s}}\leq\left(\frac{7}{5}\right)^{m-s}.$
Proof 15
We first show that the function $f(x)=\frac{x\ln(a^{2})+\ln(x)}{(7/5)^{x}}$ is strictly decreasing for $x\geq 5$, where $a=\max\{T^{\frac{1}{2({\mathcal{M}}+1)}},2\}\geq 2$. Notice that the derivative function
$<$ $\,\frac{\ln(m+4)+(m+4)\ln(a^{2})}{\ln(s+4)+(s+4)\ln(a^{2})}$
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$.
Lemma 17
For any epoch $m$, the following inequality holds:
$\sum_{s=1}^{m}\lambda_{s}\leq 2a^{8}\ln(a)(m^{2}+3m(2+\ln(K))).$
Proof 16
Since $a\geq 2$ and $K\geq 2$, we can have
$=\sum_{s=1}^{m}a^{8}(2(s+4)\ln(a)+\ln(s+4)+\ln(4K^{2}\ln(aK))$
$\leq\sum_{s=1}^{m}a^{8}(2(s+4)\ln(a)+2(s+1)\ln(a)+6\ln(K)\ln(a))$
$\leq\sum_{s=1}^{m}a^{8}\ln(a)(4s+10+6\ln(K))$
$=2a^{8}\ln(a)(m^{2}+3m(2+\ln(K))).$
Lemma 18
Proof 17
By Lemma 5, we can get
A union bound over the $K$ arms conclude the proof.
Since $m\geq 1$, $K\geq 2$ and $a\geq 2$, then we can get
$N_{m}$ $=K\lambda_{m}a^{2(m-1)}$
$=Ka^{2(m+3)}\ln(4K^{2}(m+4)a^{2(m+4)}\ln(aK))$
$\leq Ka^{2(m+3)}(2(m+4)\ln(a)+\ln(m+4)+\ln(K^{2}\ln(aK))+\ln(4))$
$\leq Ka^{2(m+3)}(2(m+4)\ln(aK)+(m+4)\ln(aK)+3\ln(aK)+2\ln(aK))$
$\leq Ka^{2(m+3)}((4m+16)\ln(aK))$
$=Ka^{2(m+4)}((m+4)\ln(aK))$
$=1/\delta_{m}.$
This way we can always guarantee the following inequality:
$\rho_{m}:=\sum_{s=1}^{m}\frac{2^{m-s}D_{s}}{a^{4(m-s)}N_{s}}.$
Lemma 19
For all epochs $m$ and arms $k$, we can have
Proof 18
$-\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
Additionally, since
we can get
$-\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
We now establish the upper bound for $\Delta_{k}^{m}$ by induction on $m$.
Assuming the statement is true for the case of $m-1$, we have
Where the second inequality follows from the induction hypothesis.
Next, we provide the lower bound of $\Delta_{k}^{m}$. We can get
D.3 Proving Theorem 4
We first define the regret $R_{k}^{m}$ generated by arm $k$ in epoch $m$ as
Then we analyze the regret in following three cases:
Case 1:
$0<\Delta_{k}\leq 64\rho_{m-1}$.
If $k\neq k_{m}$, then we have
$R_{k}^{m}=\Delta_{k}n_{k}^{m}\leq 64\rho_{m-1}n_{k}^{m}.$
If $k=k_{m}$, then we have
$R_{k}^{m}=\Delta_{k}\widetilde{n}_{k_{m}}^{m}\leq 64\rho_{m-1}N_{m}.$
Case 2:
$\Delta_{k}\leq 4\cdot a^{-(m-1)}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
If $k\neq k_{m}$, then we have
If $k=k_{m}$, since $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$, we can get:
Case 3:
$\Delta_{k}>4\cdot a^{-(m-1)}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
By Lemma 7 we have
[…]
$\mathbb{E}[D_{m}]=2(1-\delta_{m})C_{m}+\delta_{m}N_{m}\leq 2C_{m}+1$
and we can bound $\sum_{m=1}^{M}\rho_{m-1}N_{m}$ as
Combining Eq. (10) and Eq. (11), by Lemma 17, since $a=\max\{T^{\frac{1}{2(M+3)}},2\}$ and $M=\min\{{\mathcal{M}},\log_{2}(T)\}$, we can get
$\begin{split}{\mathbb{E}}[R(T)]&\leq\sum_{m=1}^{M-1}570a^{2}{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}\sum_{m=1}^{M-1}1140a^{2}C_{m}+570a^{2}+2^{5}a^{8}\ln(a)\left(\sum_{\Delta_{k}>0}\frac{M^{2}+3M(2+\log(K))}{\Delta_{k}}\right.\\ &\quad\left.+\frac{K(\log^{2}(8/\Delta)+3\log(8/\Delta)(2+\log(K))}{\Delta}\right)+\log_{2}(8/\Delta)\\ &=O\left(CT^{\frac{1}{M+3}}+T^{\frac{4}{M+3}}\left(\sum_{\Delta_{k}>0}\frac{M\log(KT)}{\Delta_{k}}+\frac{K\log(T)\log(1/\Delta)\log(K/\Delta)}{M\Delta}\right)\right).\end{split}$
D.4 Proving for Theorem 5
Proof 19
In the context of batched bandits, let the static grids be defined as $0=t_{0}<t_{1}<\cdots<t_{{\mathcal{M}}}=T$. According to Lemma 2 in Gao et al. [2019], the lower bounds for stochastic batched bandits are given by:
$R(C=0,{\mathcal{M}})=O\left(K\cdot\max_{j\in[{\mathcal{M}}]}\frac{t_{j}}{t_{j-1}+1}\right).$ (12)
For adversarial corruption, we define an adversary who starts attacking at round 00 and continues until the end of a batch, denoted by $t_{s}$ for $0<s<{\mathcal{M}}$. The attack strategy involves setting the rewards of all arms to zero, forcing any algorithm to pull all arms evenly in round $s+1$, which leads to a regret of approximately $t_{s}-t_{s-1}$. Therefore, under adversarial corruption, the lower bound can be expressed as:
$\begin{split}R(C,{\mathcal{M}})&=O\left(K\cdot\max_{j\in[{\mathcal{M}}]}\frac{t_{j}}{t_{j-1}+1}+C\cdot\max_{j\in[{\mathcal{M}}-1]}\frac{t_{j+1}-t_{j}}{t_{j}}\right)\\ &=O\left(Kt_{1}+(K+C)\max_{j\in[{\mathcal{M}}-1]}\frac{t_{j+1}-t_{j}}{t_{j}}\right)\\ &=O\left(Kt_{1}+(K+C)\left(\frac{T}{t_{j}}\right)^{\frac{1}{{\mathcal{M}}-1}}\right).\end{split}$ (13)
According the formula, we can get the optimal value of $t_{1}=T^{\frac{1}{{\mathcal{M}}}}\left(\frac{K({\mathcal{M}}-1)}{K+C}\right)^{\frac{{\mathcal{M}}-1}{{\mathcal{M}}}}$. Combine equation 13, we can get
$R(C,{\mathcal{M}})\geq\Omega\left(T^{\frac{1}{{\mathcal{M}}}}\left(K+C^{1-\frac{1}{{\mathcal{M}}}}\right)\right).$
Appendix E Proof of Theorem 6
E.1 Notations
E.2 Lemmas for Proving Theorem 6
Lemma 20
For the MA-BARBAT algorithm time horizon $T$, the number of epochs $M$ is at most $\log(VT)$. In the $m$-th epoch, the selected arm $k_{m}$ must satisfy $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ for all agents $v\in[V]$.
Proof 20
The length of epoch $m$ is given by $N_{m}=\lceil K\lambda_{m}2^{2(m-1)}\rceil\geq 2^{2(m-1)}/V$. From the lower bound of $N_{m}$, we can complete the first statement. Since $\Delta_{k}^{m-1}\leftarrow\max\{2^{-(m-1)},r_{*}^{m}-r_{k}^{m}\}$, there exists at least one arm that satisfies $\Delta_{k}^{m-1}=2^{-(m-1)}$ and all arms satisfy $\Delta_{k}^{m-1}\leq 2^{-(m-1)}$. Since $r_{k_{m}}^{m-1}>r_{*}^{m-1}$, the equality $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$ must hold.
Lemma 21
For epoch $s$ and $m$ with $1\leq s\leq m$, the following inequality holds:
$\frac{\lambda_{m}}{\lambda_{s}}\leq\left(\frac{7}{6}\right)^{m-s}.$
Proof 21
where we use the monotonicity of $f(x)$ and the fact that $K\geq 2$ and $V\geq 1$.
Lemma 22
For any epoch $m$, the following inequality holds:
$\sum_{s=1}^{m}\lambda_{s}\leq 2^{8}(m^{2}+m(10+\ln(VK))).$
Proof 22
$=\sum_{s=1}^{m}2^{8}(2(s+4)\ln(2)+\ln(s+4)+\ln(4VK^{2}\ln(VK))$
$\leq\sum_{s=1}^{m}2^{8}(2s+7+\ln(4V^{2}K^{3}))$
$=2^{8}(m^{2}+8m+m\ln(4V^{2}K^{3}))$
$\leq 2^{8}(m^{2}+m(10+3\ln(VK))).$
Lemma 23
Proof 23
We can divide the sum to two components:
By applying the Chernoff-Hoeffding inequality (Theorem 1.1 in Dubhashi and Panconesi [2009]), we can get
By applying the concentration inequality for martingales (Theorem 1 in Beygelzimer et al. [2011]), with probability at least $1-\frac{\beta_{m}}{4}$, we have
Since $q_{k}^{m}=\widetilde{n}_{k}^{m}/N_{m}$, $\sum_{t\in E_{m}}\sum_{v\in[V]}|C_{v,k}^{t}|\leq C_{m}$ and $n_{k}^{m}\geq\lambda_{m}\geq 16\ln(4/\beta_{m})/V$, with probability at least least $1-\frac{\beta_{m}}{4}$, we can get the following inequality:
Similarly, $-\frac{B_{k}^{m}}{V\widetilde{n}_{k}^{m}}$ also meets this bound with probability $41-\beta/$ 4. Therefore, we have
Combine Eq. 14 and Eq. 15, we complete the proof.
We also define an event ${\mathcal{E}}_{m}$ for epoch $m$ as follows:
Lemma 24
Proof 24
By Lemma 23, we can get
A union bound over the $V$ agents and the $K$ arms conclude the proof.
Since $m\geq 1$, $V\geq 1$ and $K\geq 2$, then we can get
$VN_{m}$ $=KV2^{2(m-1)}\lambda_{m}$
$=K2^{2(m+3)}\ln((m+4)2^{2(m+5)}VK^{2}\ln(K))$
$\leq K2^{2(m+3)}(\ln(m+4)+2(m+5)\ln(2)+\ln(VK^{3}))$
$\leq K2^{2(m+4)}((m+4)\ln(VK))$
$\leq 1/\delta_{m}.$
$D_{m}=\begin{cases}2C_{m}&\textit{when ${\mathcal{E}}_{m}$ occurs}\\ VN_{m}&\textit{when ${\mathcal{E}}_{m}$ dose not occur}\end{cases}.$
This way we can always guarantee the following inequality:
By the definition of $D_{m}$, we have
$\rho_{m}:=\sum_{s=1}^{m}\frac{D_{s}}{8^{m-s}VN_{s}}.$
Then we have the following lemma.
Lemma 25
For all epochs $m$ and arms $k$, we can have
Proof 25
$-\frac{D_{m}}{VN_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{VN_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}.$
Additionally, since
we can get
$-\frac{D_{m}}{VN_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k^{*}}^{m}}}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{VN_{m}}.$
We now establish the upper bound for $\Delta_{k}^{m}$ by induction on $m$.
Assuming the statement is true for the case of $m-1$, we have
Where the second inequality follows from the induction hypothesis.
Next, we provide the lower bound of $\Delta_{k}^{m}$. We can get
E.3 Proof for Theorem 6
Then we analyze the regret in following three cases:
Case 1:
$0<\Delta_{k}\leq 64\rho_{m-1}$.
If $k\neq k_{m}$, then we have
$R_{v,k}^{m}=\Delta_{k}n_{k}^{m}\leq 64\rho_{m-1}n_{k}^{m}.$
If $k=k_{m}$, then we have
$R_{v,k}^{m}=\Delta_{k}\widetilde{n}_{k_{m}}^{m}\leq 64\rho_{m-1}N_{m}.$
Case 2:
$\Delta_{k}\leq 8\cdot 2^{-m}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
If $k\neq k_{m}$, then we have
If $k=k_{m}$, since $\Delta_{k_{m}}^{m-1}=2^{-(m-1)}$, we can get:
Case 3:
$\Delta_{k}>8\cdot 2^{-m}$ and $\rho_{m-1}\leq\frac{\Delta_{k}}{64}$.
By Lemma 7 we have
[…]
$\mathbb{E}[D_{m}]=2(1-\delta_{m})C_{m}+\delta_{m}N_{m}\leq 2C_{m}+1$
and we can bound $\sum_{m=1}^{M}\rho_{m-1}N_{m}$ as
Combining Eq. (16) and Eq. (17), by Lemma 22, we can get
$\begin{split}{\mathbb{E}}[R_{i}(T)]&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(VT)+3\log(VT)\log(30VK))}{V\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30VK)))}{V\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880\frac{C_{m}}{V}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(VT)\log(30VKT)}{V\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240VK/\Delta)}{V\Delta}+\log(8/\Delta)\\ &=O\left(\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log(T)\log(VKT)}{V\Delta_{k}}+\frac{K\log(1/\Delta)\log(VK/\Delta)}{V\Delta}\right).\end{split}$
So the cumulative regret for all agents as
${\mathbb{E}}[R(T)]=O\left(C+\sum_{\Delta_{k}>0}\frac{\log(VT)\log(VKT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(VK/\Delta)}{\Delta}\right).$
$\textrm{Cost}(T)=\sum_{v\in[V]}M=O(V\log(VT)).$

### v2

@@ B.5.3 Proof for Theorem 6
We define ${\mathcal{A}}^{m}\triangleq\left\{k\in[K]\,\big|\,0<\Delta_{k}\leq 64\rho_{m-1}\right\}$ for epoch $m$. By combining all three cases, we can upper bound the regret as
@@
$\begin{split}R(T)&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[dD_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16Kd\lambda_{m}}{\Delta}+d\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[dD_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}Kd(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+d\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880dC_{m}+1440d+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}Kd\log(8/\Delta)\log(240K/\Delta)}{\Delta}+d\log(8/\Delta)\\ &=O\left(dC+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{dK\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
Appendix C Experimental Details
C.1 Cooperative Multi-Agent Multi-Armed Bandits
We first consider CMA2B, where the mean rewards $\{\mu_{k}\}_{k\in[K]}$ is uniformly distributed in the interval $[0.02,0.96]$. For each arm $k$, we generate the stochastic rewards according to a truncated normal distribution with support $[0,1]$, mean $\mu_{k}$, and variance $0.1$. We use a time horizon $T=50000$, set the corruption level $C=2000,5000$ and the number of arms $K=12,16$. Following  [lu2021stochastic], we set the two worst arms as target arms and corrupt the target arm’s reward to $1$ and the other arms’ rewards to zero until the corruption level is exhausted.
As seen in Table 2, the time cost by BARBAT is much shorter than that of the FTRL, which confirms that the computational efficiency of the BARBAT framework is much higher than that of the FTRL framework.
BARBAT DRAA IND-BARBAR IND-FTRL
K = 12 0.13 s 0.14 s 0.12 s 1.53 s
K = 16 0.13 s 0.15 s 0.14 s 1.91 s
Table 2: The cost time of each agent for BARBAT and all baseline methods.
Refer to caption Refer to caption
(a) K = 12 (b) K = 16
Figure 4: The feedback structure for the strongly observable graph bandits.
BARBAT HYBRID LBINF LBINF_LS LBINF_GD
K = 12 0.36 s 14.63 s 10.16 s 10.21 s 9.84 s
K = 16 0.37 s 18.62 s 12.37 s 12.67 s 12.69 s
Table 3: The cost time of each agent for BARBAT and all baseline methods.
C.2 Strongly Observable Graph Bandits
Following [lu2021stochastic], we adopt the Erdos–Renyi model [erd6s1960evolution] to generate feedback graph in Figure 4, with the independence number $\alpha=4$. Specifically, for each pair of arms $(u,v)\in[K]\times[K]$ with $u=v$, we connect them with a fixed probability $0.5$. Blue circles represent vertices without self-loops, while white circles represent vertices with self-loops.
C.3 $d$-Set Semi-bandits
In $d$-set semi-bandits, we set $d=3$ for $K=12$, and $d=4$ for $K=16$. Following the recent works [zimmert2019beating, tsuchiya2023further], we generate the stochastic rewards according to a Bernoulli distribution, other settings are the same as cooperative standard MAB.
As seen in Table 3, the time cost by BARBAT is much shorter than that of the FTRL, which confirms that the computational efficiency of the BARBAT framework is much higher than that of the FTRL framework.